OpenStudy (anonymous):

Please help: write the polynomial in completely factored form f(x)=x^3-x-6

OpenStudy (lieutenantgeneral):

x = 2 (x - 2) (x^2 + 2x + 3)

OpenStudy (anonymous):

Thanks. How does x=2?

OpenStudy (anonymous):

@LieutenantGeneral

OpenStudy (lieutenantgeneral):

2 * 3 = 6 - 3 - 6 go to wolframalpha.com it'll show everything

OpenStudy (anonymous):

ok thank you

OpenStudy (lieutenantgeneral):

Welcome

OpenStudy (anonymous):



OpenStudy (anonymous):

thank you @surjithayer :)

OpenStudy (anonymous):

yw